OpenStudy (anonymous):

The probability of a student scoring 75% in class work is 0.64, and the probability of a student scoring 85% is 0.45. Event A: The student scores 75%. Event B: The student scores 85%. The probability of a student scoring 85% in class work, given that they have already scored 75% in class work, is 0.55. The probability of a student scoring 75% in class work, given that they have already scored 85% in class work, is 1. Which statement is true?

OpenStudy (anonymous):

Events A and B are independent because P(A|B) = P(A). Events A and B are independent because P(A|B) = P(B). Events A and B are independent because P(A|B) = P(A) + P(B). Events A and B are not independent because P(A| B) ≠ P(A). Events A and B are not independent because P(A|B) = P(A).

OpenStudy (amistre64):

your material should define independence for you. what is that definition?

OpenStudy (anonymous):

Oh I forgot to say that I think it's E

OpenStudy (amistre64):

if P(A|B) = P(A|U) then we have independance since the amount of A in B is the same ratio as the amount of A in the universal set

OpenStudy (anonymous):

So I'm worng

OpenStudy (amistre64):

The probability of a student scoring 75% in class work is 0.64, P(A|U), pr just P(A) and the probability of a student scoring 85% is 0.45. P(B|U), or simply P(B) Event A: The student scores 75%. Event B: The student scores 85%. The probability of a student scoring 85% in class work, given that they have already scored 75% in class work, is 0.55. P(A|B) The probability of a student scoring 75% in class work, given that they have already scored 85% in class work, is 1. P(B|A)

OpenStudy (amistre64):

i see no independence for anything ... P(A|B) not equal P(A) P(B|A) not equal P(B)